# StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models

### v2

@@ Abstract
Planning and acting to solve ‘real’ tasks using large language models (LLMs) in interactive environments has become a new frontier for AI methods. While recent advances allowed LLMs to interact with online tools and solve robotics tasks, long range reasoning tasks remain a problem for LLMs. Existing methods to address this issue are very resource-intensive and require additional data or human crafted rules; instead, we propose StateAct, a simple method based on few-shot in-context learning alone. StateAct enhances ‘chain-of-thought’ with state- and goal-tracking to help with planning and acting with LLMs. The method establishes the new state-of-the-art on Alfworld for in-context learning methods (+14% over the previous best few-shot in-context learning method) and performs on par with methods that use additional training data and additional tools such as code-execution. We also demonstrate that StateAct with enhanced ‘chain-of-states’ allows the agent to both solve longer horizon problems and to be more efficient in the number of steps required to solve a task. We show that StateAct works across a variety of API-based and open source LLMs. Finally, we also conduct ablation studies and show that ‘chain-of-thoughts’ help state-tracking accuracy, while a json-structure harms overall performance. We open-source our code and annotations at https://github.com/ai-nikolai/stateact.
StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models
@@ 1 Introduction
Using the in-built world and commonsense knowledge [^1] Commonsense- and world- knowledge as explored by Lauscher et al. (2020), for example. of large language models (LLMs), such as GPT-3, Mixtral and Gemini Brown et al. (2020); Jiang et al. (2024); Anil et al. (2023) to perform interactive reasoning tasks has become a frontier in AI research, with “AI Agents” now able to solve a range of multi-modal complex tasks Durante et al. (2024). These range from solving simulated robotics tasks Puig et al. (2018); Shridhar et al. (2021) and digital tasks, such as online shopping Yao et al. (2023a) and navigating operating systems Liu et al. (2023), to playing a variety of games Côté et al. (2019); Liu et al. (2023).
While LLMs are increasingly successful in solving AI agent tasks, Li et al. (2023b) show that LLMs struggle with long range understanding and Coelho et al. (2024) show that LLMs often focus on the beginning of the textual history, further amplifying the problem. Existing efforts to address these issues are resource-intensive: Wu et al. (2024) required human expert annotations of rules, Sun et al. (2023) required a code execution environment with carefully crafted code-based prompts, and Fu et al. (2024) used additional training data together with retrieval augmented generation (RAG) to help the AI agent.
To overcome these challenges, we introduce “StateAct”, a novel LLM-agent based on few-shot in-context learning. The method tracks the goal by ‘reminding’ the agent of the goal. StateAct, also, explicitly keeps track of the agent’s state (such as location and inventory). Additionally, we utilise ‘chain-of-thought’ (i.e. explicit reasoning Wei et al. (2023)) as an optional addition. At the core of the method lies a straightforward annotation of the few-shot examples, specifically with goal-, state- and reasoning- traces, which we call ‘chain-of-states’.
Our method establishes a new state-of-the-art for Alfworld Shridhar et al. (2021) for few-shot in-context learning-based methods, outperforming the previous best method by 14% and even outperforming methods that use external tools by 2.5%. Furthermore, while our method does not require additional data or external tools and only minimal and easy human annotations, we perform on par with the current state-of-the-art that is much more resource-intensive and requires additional data and expert human annotations. Comparing to in-context learning methods, we also demonstrate that our method is more efficient in terms of the number of steps taken to solve a task and is able to solve longer-horizon tasks, validating the hypothesis that explicit state-tracking and goal-tracking help with solving longer horizon reasoning and planning tasks.
@@ 2 Background
AI agents have historically used reinforcement learning (RL) to solve tasks Sutton and Barto (2018). With the capabilities of LLMs improving, recent work has combined LLMs and RL, training additional policies or value functions to make predictions (Li et al., 2022; Nottingham et al., 2023).
2.1 In-context learning approaches
Huang et al. (2022a, b) and Singh et al. (2022) were among the first to use LLMs directly to act in an interactive environment; their method produces agent actions as output after receiving environment observations as input.
ReAct Yao et al. (2023b) took this work further by combining ‘acting‘ Huang et al. (2022a) and ‘chain-of-thought‘ Wei et al. (2023). ReAct establishes state-of-the-art for for approaches based only on in-context learning and while it is a very scalable method, the performance (i.e. success rate) is still limited.
ExpeL Zhao et al. (2023) uses additional training data to generate ‘insights’ and ‘success trajectories’ during training. At inference time, they look up the closest ‘success trajectories’ as few-shot examples to the agent (as opposed to the fixed few-shot examples that we use) and augment them with these ‘insights’ to perform the final inference. They achieve 59% on Alfworld using retrieved ‘success trajectories’ (as few-shot examples) + ‘insights’ and 50% using the same few-shot examples as ReAct + ‘insights’.
2.1.1 In-context learning and additional tools
@@
The current state-of-the-art for in-context learning-based approaches in combination with additional tools is ProgPrompt (Singh et al., 2022) and follow-on work AdaPlanner Sun et al. (2023). AdaPlanner introduces a code-based prompt Li et al. (2023a). They use code-execution as an additional tool, by executing LLM generated code and feeding the results into the next LLM generation. The shortcoming of AdaPlanner is that it requires very complex human-crafted prompts that are hard to scale to new environments and it requires the additional step of code-execution.
@@ 2.2 State tracking in LLM-based agents
Concurrent work to ours, AutoGuide Fu et al. (2024), uses ReAct as the base agent as well as additional training data to create ‘state-aware’ text-based guidelines for the LLM-agent; they then use a type of retrieval augmented generation (RAG) process to guide the decision making process. They embed the current observation as a summary (which they call ‘state-aware’) and use an LLM to ‘look up’ the relevant ‘state-aware’ guideline, which is then fed into a final LLM to generate an action. Using training data and LLM-based RAG, they achieve 79.1% on top of a ReAct agent on Alfworld. Their training and RAG approach could be used complementary to our StateAct LLM-agent [^2] However, the cost of using RAG at every generation step could be significantly higher than our method..
Chen et al. (2024) propose state-tracking as a way to help the agent solve the task without training data. Their method differs to ours two-fold. Firstly, they employ a complex sequence of components working together (an LLM-based attention over the observations, an LLM-based compilation of a complex state and, finally, a prediction of a program). Secondly, their system involves execution of actual programs. Our method, on the other hand, requires a straightforward extension of ‘chain-of-thought’ and uses a single LLM call to produce the state, thought and action. Additionally, we do not require program execution.
2.2.1 Fine-tuning approaches
Previous fine-tuning approaches did not significantly enhance performance Zhou et al. (2024); Yao et al. (2023b); Shridhar et al. (2021). A concurrent work, however, ActRe Yang et al. (2024b) achieves 83% by fine-tuning gpt-3.5-1106 on additional training data.
2.3 “Multi-Agent Conversation” approaches
A new trend is to use multiple LLMs concurrently to ‘chat’ to one another to produce a result. A recent work in this direction by Wu et al. (2023) achieves 67% on Alfworld.
2.4 Joined rule and LLM-based agents
StateFlow Wu et al. (2024) uses Finite State Machines (FSMs) combined with LLMs to solve Alfworld. These FSMs are human expert-crafted states, transitions and rule-based heuristics, where the LLM is asked to perform limited tasks in each of the given states. While their performance is 82% on Alfworld, we believe this result is more comparable to a rule-based ‘expert’ than an LLM-based agent (notably close to 100% can be achieved on Alfworld using hand-crafted ‘experts’ Shridhar et al. (2021)).
2.5 Approaches with multiple trials
Works such as Reflexion Shinn et al. (2023) or work by Sun et al. (2023); Yang et al. (2024a) are methods that perform several evaluations on the same test environments and perform learning loops (during evaluation) on these test-set environments to improve the score. While the score necessarily goes up (as only the failed environments are re-run), we do not dive into this as it does not compare to the actual agent performance; additionally, ‘reflection‘ can be used on top of most LLM-based agents (e.g. Reflexion uses ReAct as the base agent).
@@ 3 Method
Figure 1: StateAct interacts with the environment directly, without additional tools or resources (a). A skeleton of the StateAct prompt (b).
StateAct is an LLM-based AI agent that works on top of pre-trained large language models. StateAct takes the textual ‘observation’ from the environment and, after a single call to the pre-trained LLM, returns the ‘action’ back to the environment (without the use of additional tools or resources), see Figure 1(a).
StateAct utilises in-context learning Brown et al. (2020); Wei et al. (2023) to make the agent interact with the environment. At the core of the approach is a prompt that consists of few-shot examples of successful ‘interaction traces’ as well as the current interaction trace up to the current step in the environment, see Figure 1(b). An interaction trace consists of alternating observations from the environment and desired (or actual) outputs from the LLM. In the case of StateAct, the LLM is tasked to generate the ‘goal’, ‘state’, ‘thought’ and ‘action’. The action is then extracted and passed to the environment to produce the next observation, see Figure 2. For example, a sample of the trace for step n in Alfworld is:
@@
Current Location: countertop 1
@@
Thought: I find the tomato, now I need to take it.
@@
In the above example, the agent produces the action ‘take tomato 1’ in response to the observation ‘On the countertop 1, you see a tomato 1’.
Figure 2: An example of how StateAct is working. The depicted traces are from the Alfworld environment.
3.1 Goal-reminders and state-tracking using LLMs
In order to make StateAct more precise, we introduce goal and state tracking. We denote $\pi$ as the policy of an AI-agent, in the standard case at time step $t$, the policy predicts action $a_{t}$, given the history of observations and actions $[o_{t},a_{t-1},...,a_{0},o_{0}]$.
$\pi(a_{t}|o_{t},a_{t-1},...,a_{0},o_{0})$ (1)
For our case, we need to enhance the policy to incorporate the ‘state’. Similar to previous work Yao et al. (2023b), we introduce the context vector, $c_{t}$. The context vector contains the action, as well as the additional predictions of the agent , i.e. $c_{t}=(g_{0},s_{t},r_{t},a_{t})$. Where $g_{0}$ is the goal and it always remains the same (for a given environment) and uses the goal extracted from $o_{0}$; $s_{t}$ represents the predicted state at time step $t$; $r_{t}$ represents ‘chain-of-thought’-style ‘reasoning’ at time step $t$; and $a_{t}$ represents the action at time step $t$, as before. The new policy $\pi$ then becomes:
$\pi_{contextual}(c_{t}|o_{t},c_{t-1},...,c_{0},o_{0})$ (2)
@@ 4 Experimental setup
Our aim is to study the long-range acting, planning and reasoning capabilities of LLM-based agents. To achieve this, in line with previous work, we turn to simulated environments as an evaluation framework and to API-based state-of-the-art large language models. Specifically, we use Alfworld Shridhar et al. (2021), a household robotics environment, and Webshop Yao et al. (2023a), an online shopping environment, as our simulated environments. In line with previous work, we use OpenAI’s [^3] https://openai.com, last accessed June 2024. gpt-3.5-turbo-1106. We, also, evaluate Alfworld on additional LLMs and validate our results.
@@ 4.1 Alfworld
Alfworld Shridhar et al. (2021) is based on a 3D, visual, household robotics environment called Alfred Shridhar et al. (2020), which was translated into a text-based environment for ease of use for language-based AI models, see Figure 3. Alfworld has a total of 134 test-set examples and six environment types. It features long-time horizons, partial observability, an out-of-distribution evaluation set and text-based interactions. Alfworld simulates a household environment with a household assistant robot tasked with solving problems, e.g. clean an apple and put it on a table. The robot (or agent) then needs to perform a series of ‘high-level’ operations to accomplish the tasks, e.g. ‘go to fridge 1’, ‘open fridge 1’. At every step, the environment provides either a textual observation or feedback that the command has failed, e.g. ‘You open the fridge 1’, ‘You see apple 1’. The underlying text engine is based on Textworld Côté et al. (2019), see Appendix A for a complete list of the commands and further details on the environments.
4.1.1 Alfworld correction
Methods such as AdaPlanner Sun et al. (2023) have avoided this issue because they use code-based prompts and regex parsers. However, methods such as ReAct Yao et al. (2023b) and ExpeL Zhao et al. (2023) have been affected by it therefore lowering their potential performance. We report results for ReAct and StateAct both with and without these corrections.
@@ 4.2 Webshop
Webshop Yao et al. (2023a) is a simulation of an online shopping experience. Given a task, e.g. “I want a blue water-proof winter jacket, less than \$100”, the agent needs to search a product catalogue, browse through the search results, select the most fitting product, select the attributes, e.g. colour, size, and then buy the product. In line with previous work, we use the text-based version of Webshop, where all descriptions of the website are given in text form, see Figure 4. Webshop features a realistic large-scale product catalogue, a search engine and very varied product attributes depending on the category of product, see Appendix B for more details. In total, the test set consists of 500 examples and each one is of the type “search and buy a product”. Overall, Webshop has a maximum of 15 steps and two commands: 1. search[<query>], 2. click[<button>].
4.3 In context learning
Since ReAct Yao et al. (2023b) forms the underlying agent for many current Zhao et al. (2023) and state-of-the-art approaches Fu et al. (2024), we use the same few-shot ‘interaction traces’ as ReAct. The main reason is to have a fair comparison and to isolate additional effects - such as performance change - from different in-context examples.
In total, Alfworld has six types of tasks and ReAct uses two in-context examples per task type to prompt the language models. On average, each ReAct example ranges from 352 words to 591 words (590 tokens to 935 tokens). For our study, we reuse the observations, thoughts and actions, and annotate these examples further with goal ‘reminders’ and state tracking, which results in a range from 484 to 911 words (807 tokens to 1458 tokens) per example. During our annotation, we discovered minor errors in the ReAct prompts and fixed them as well. We release all our annotations with our code release. In comparison, AdaPlanner Sun et al. (2023), uses a different code-based approach and the prompt has 1104 words (2015 tokens) on average.
We use the two-shot examples from ReAct for Alfworld and the one-shot example from ReAct for Webshop in all our experiments.

### v3

# StateAct: Enhancing LLM Base Agents via Self-prompting and State-tracking

@@ Abstract
Large language models (LLMs) are increasingly used as autonomous agents, tackling tasks from robotics to web navigation. Their performance depends on the underlying base agent. Existing methods, however, struggle with long-context reasoning and goal adherence. We introduce StateAct, a novel and efficient base agent that enhances decision-making through (1) self-prompting, which reinforces task goals at every step, and (2) chain-of-states, an extension of chain-of-thought that tracks state information over time. StateAct outperforms ReAct, the previous best base agent, by over 10% on Alfworld, 30% on Textcraft, and 7% on Webshop across multiple frontier LLMs. We also demonstrate that StateAct can be used as a drop-in replacement for ReAct with advanced LLM agent methods such as test-time scaling, yielding an additional 12% gain on Textcraft. By improving efficiency and long-range reasoning without requiring additional training or retrieval, StateAct provides a scalable foundation for LLM agents. We open source our code to support further research at https://github.com/ai-nikolai/stateact.
StateAct: Enhancing LLM Base Agents via Self-prompting and State-tracking
@@ 1 Introduction
Leveraging the in-built world and commonsense knowledge [^1] Commonsense- and world- knowledge as explored by Lauscher et al. (2020), for example. of large language models (LLMs), such as GPT, Gemini and Mixtral Brown et al. (2020); Anil et al. (2023); Jiang et al. (2024) to perform interactive reasoning tasks has become a frontier in AI research. “AI Agents” are now able to solve a range of multi-modal complex tasks Durante et al. (2024).
These include simulated robotics tasks Puig et al. (2018); Shridhar et al. (2021) and digital tasks, such as online shopping Yao et al. (2023a), navigating operating systems Liu et al. (2023), and playing a variety of games Côté et al. (2019); Liu et al. (2023); Prasad et al. (2024).
At the core of an LLM agent is the base agent, such as Act (Huang et al., 2022a), ReAct (Yao et al., 2023c), and AdaPlanner (Sun et al., 2023). Existing efforts to improve LLM agents build on top of base agents and are usually quite resource-intensive: Wu et al. (2024) require human expert annotations of rules; Sun et al. (2023) require a code execution environment with carefully crafted code-based prompts; Fu et al. (2024) use additional training data together with retrieval augmented generation (RAG) to help the AI agent; Yang et al. (2024) use additional training data to fine-tune the LLM; Shinn et al. (2023) and Prasad et al. (2024) use test-time computation to produce better results. Notably, most of the current state-of-the-art methods use ReAct as the base agent.
In our work, we propose a new base agent, called StateAct. Starting with the observation that: i) LLM agents fail to follow the original instruction and goal in longer interactions; and ii) LLMs struggle with long-context despite longer available contexts (Li et al., 2023b; Coelho et al., 2024), we propose two main contributions for improving the base agent. To address i), we propose a mechanism for the agent to ‘self-prompt’ at every turn of the interaction to improve staying on track with the main goal. Our agent ‘reminds’ itself of the goal at every turn. To address ii), we propose ‘chain-of-states’, an extension of chain-of-thought based on state tracking to help the agent stay on track with the current interaction and context. The agent keeps track of its state in the environment (such as location and inventory).
Our experiments show that StateAct achieves near state-of-the-art performance without using additional training data or additional tools, and significantly outperforms ReAct across multiple tasks and eight frontier LLMs of varying sizes. Specifically, StateAct improves performance over ReAct by more than 10% on Alfworld (Shridhar et al., 2021), 30% on Textcraft (Prasad et al., 2024), and 7% on Webshop (Yao et al., 2023a).
Additionally, we validate that StateAct can serve as a drop-in replacement for existing extension methods. Using test-time computation (Prasad et al., 2024), we achieve a further 12% performance gain with StateAct on Textcraft using ADaPT.
@@ 2 Background
Figure 1: Overview of LLM-agent methods and their components.
Due to recent advances, LLMs are now being used as autonomous agents in interactive environments as an alternative to traditional reinforcement learning (RL) Sutton and Barto (2018); Yao et al. (2023c); Li et al. (2022); Nottingham et al. (2023). LLM agents now tackle tasks in simulated environments such as Alfworld, Webshop, and Textcraft. LLM agents consist of the base agent and a possible extensions, such as fine-tuning (FT), retrieval augmented generation (RAG), test-time scaling (TTS), repeated attempts (REPS) or tools; see Table 1. StateAct is a base agent that can be used in combination with extensions.
2.1 Base agents
Huang et al. (2022a, b) were among the first to use LLMs directly to act in an interactive environment; their method produces agent actions as output after receiving environment observations as input. ReAct Yao et al. (2023c) take this work further by combining acting Huang et al. (2022a) and chain-of-thought Wei et al. (2023). Progprompt and AdaPlanner (Singh et al., 2022; Sun et al., 2023) use code-based prompts to interact with environments. Notably, ReAct is the base agent of modern state-of-the-art approaches. Our method, StateAct, falls in the category of base agents and is therefore most comparable to ReAct.
2.2 Extensions of base agents: RAG, fine-tuning, test-time-scaling
ExpeL Zhao et al. (2023) extend ReAct by using additional training data to generate success trajectories during training. At inference time, they look up the closest success trajectories as few-shot examples to the agent. Follow-on work, AutoGuide Fu et al. (2024), uses ReAct as the base agent with additional training data to create state-aware text-based guidelines. This ‘knowledge’ of the environment is then used with retrieval augmented generation (RAG) to guide the decision-making process. While AutoGuide achieves the best result among RAG approaches, the complexity of the setup makes it less scalable in practice.
Chen et al. (2023); Yao et al. (2023c) introduce fine-tuning of ReAct with marginal improvements. KnowAgent (Zhu et al., 2025) compile knowledge of the environment and distill this into the LLM to produce better results. The best approach, ActRe Yang et al. (2024), achieves successful fine-tuning of ReAct by annotating successful trajectories with CoT before fine-tuning. Fine-tuning yields the best performance for a specific task, however, requires additional training, which is costly, and does not allow for generalisation.
ADaPT (Prasad et al., 2024) and THREAD (Schroeder et al., 2024) use test-time computation to achieve better results. This approach is in line with works such as tree-of-thought (Yao et al., 2023b), where the AI agent proposes several thoughts or actions in one go during test-time inference and another model (often the same LLM) evaluates these; the top k promising thoughts or actions are then expanded upon in a beam-search manner. These methods produce strong results and are very easy to setup, albeit they often require more compute budget compared to base agents. In our work we combine ADaPT with StateAct.
2.3 Alternative methods: tools, hand-crafted rules, multi-agent
@@
ProgPrompt (Singh et al., 2022) and follow-on work AdaPlanner (Sun et al., 2023) introduce a code-based prompts Li et al. (2023a). They use code-execution as an additional tool, by executing LLM-generated code and feeding the results into the next LLM generation. The shortcoming of such code-based prompts is that they require human experts to annotate very long prompts. This can be hard to scale to new environments and requires an additional step of code-execution.
StateFlow Wu et al. (2024) uses Finite State Machines (FSMs) combined with LLMs to solve Alfworld. These FSMs are human expert-crafted states, transitions and rule-based heuristics, where the LLM is asked to perform limited tasks in each of the given states. While this method can achieve a very high score, it is limited, as it requires a human expert to design the FSM.
Another approach is to use multiple LLMs to ‘chat’ to one another to produce a result. Multi-agent frameworks Wu et al. (2023) only achieve minor improvements over using a single agent.
@@ 2.4 State tracking in LLM-based agents
Chen et al. (2024) propose state-tracking as a way to help the agent solve the task without training data. Their method differs from StateAct two-fold. Firstly, they employ a complex sequence of components working together, which are an LLM-based attention over the observations, an LLM-based compilation of a complex state and a prediction of a program. Secondly, their system involves execution of actual programs. StateAct, on the other hand, requires a straightforward extension of CoT and uses a single LLM call to produce the state, thought and action. Additionally, it does not require program execution. STATLER (Yoneda et al., 2024) also introduce state-tracking for LLM agents. While the state has similarity to StateAct, there are notable differences. Firstly, Statler is aimed at lower level robotics execution and works with domain specific functions. Secondly, Statler produces and requires code to update and read from the state. This complex construction of the state is hard to scale to new environments.
@@ 3 Method
StateAct is an LLM-based AI agent that works on top of pre-trained LLMs. It takes the textual ‘observation’ from the environment and, after a single call to the pre-trained LLM, returns the ‘action’ back to the environment, without the use of additional tools or resources.
StateAct utilises in-context learning Brown et al. (2020); Wei et al. (2023) to make the agent interact with the environment. At the core of the approach is a prompt that consists of few-shot examples of successful interaction traces as well as the current interaction trace up to the current step in the environment. An interaction trace consists of alternating observations from the environment and desired (or actual) outputs from the LLM. In the case of StateAct, the LLM is tasked to generate the ‘goal’, ‘state’, ‘thought’ and ‘action’. The action is then extracted and passed to the environment to produce the next observation, see Figure 2. This renders StateAct similar to ReAct and therefore an easy replacement for extension methods.
3.1 Self prompting
One of the key ideas of StateAct to overcome the ‘haystack’ challenge for long prompts and therefore long horizon problems (Coelho et al., 2024) is to introduce a mechanism for the LLM to pass an instruction to itself. This approach can be applied in many settings and to alleviate various problems (including goal reminding and formatting). In our setting, we focus on keeping the Agent on track with the main goal and so the LLM reminds itself of this goal at every turn.
3.2 Chain of states as state-tracking
The idea of state-tracking is to introduce ‘structured thoughts’ into the reasoning part of the LLM Agent, specifically by giving the agent small intermediate predictions that can be inferred from the environment and actions. This method is different to existing methods such as ReAct (Yao et al., 2023c), where CoT is taken to mean verbal ‘thoughts’. The inspiration comes from the original CoT paper (Wei et al., 2023) where the LLM is tasked with producing intermediate calculations. For StateAct, the LLM agent is tasked with predicting very specific intermediate steps, such as the current location or the inventory of the agent. See Appendix K for a formalization of StateAct.
Figure 2: Example of StateAct outputs in Alfworld.
3.3 An example trace
An example of the trace for step n in Alfworld is:
@@ 3.3 An example trace
Current Location: countertop 2
@@
Locations Visited: table 1, countertop 1 Thought: I find the tomato, now I need to take it.
@@
In this example, the agent produces the action ‘take tomato 1’ in response to the observation ‘On the countertop 2, you see a tomato 1’.
@@ 4 Experimental setup
Our aim is to study reasoning capabilities of LLM-based agents. In line with previous work, we use Alfworld Shridhar et al. (2021), a household robotics environment; Webshop Yao et al. (2023a), an online shopping environment; and Texcraft (Prasad et al., 2024), a text-based tool crafting game, as our evaluation environments. We mainly evaluate StateAct as the base agent. However, we also show that StateAct can be used as a drop-in replacement for ReAct and validate it on test-time scaling. RAG (AutoGuide) and fine-tuning (ActRe)-based extensions are left for future work, as these are expensive and complex to run due to training on additional data and use of additional tools.
@@ 4.1 Alfworld
Alfworld Shridhar et al. (2021) is based on a 3D visual household robotics environment called Alfred Shridhar et al. (2020), which was translated into a text-based environment for ease of use for language-based AI models, see Figure 3. Alfworld has a total of 134 test-set examples and six environment types. It features long-time horizons, partial observability, an out-of-distribution evaluation set and text-based interactions. Alfworld simulates a household environment with a household assistant robot tasked with solving problems, e.g. clean an apple and put it on a table. The robot (or agent) then needs to perform a series of high-level operations to accomplish the tasks, e.g. ‘go to fridge 1’, ‘open fridge 1’. At every step, the environment provides either a textual observation or the feedback that the command has failed, e.g. ‘You open the fridge 1’, ‘You see apple 1’. The underlying text engine is based on Textworld Côté et al. (2019), see Appendix A for a complete list of the commands and details of the environments.
@@ 4.2 Webshop
Webshop Yao et al. (2023a) is a simulation of an online shopping experience. Given a task, e.g. “I want a blue water-proof winter jacket, less than \$100”, the agent needs to search a product catalogue, browse through the search results, select the most fitting product, select the attributes, e.g. colour, size, and then buy the product. In line with previous work, we use the text-based version of Webshop, where all descriptions of the website are given in text form, see Figure 4. Webshop features a realistic large-scale product catalogue, a search engine, and very varied product attributes depending on the category of product, see Appendix B for more details. In total, the test set consists of 100 examples and each one is of the type “search and buy a product”. Overall, Webshop has a maximum of fifteen steps and two commands: 1. search[<query>], 2. click[<button>].
4.3 Textcraft
Textcraft Prasad et al. (2024) is an environment based on the popular game Minecraft, where the task of the agent is to craft items. The environment is fully text-based. We use prompts and implementations based on the ADaPT paper (Prasad et al., 2024) that introduced this environment. We also use this environment to analyse whether StateAct performs well in combination with other methods such as ADaPT. See Appendix C for more details.
4.4 In-context learning
Since ReAct Yao et al. (2023c) forms the underlying agent for many current Zhao et al. (2023) and state-of-the-art approaches Fu et al. (2024); Yang et al. (2024); Prasad et al. (2024), we use the same few-shot interaction traces as ReAct. The main reason is to have a fair comparison and to isolate additional effects, such as performance change, from different in-context examples. Alfworld, for example, has six types of tasks and ReAct uses two in-context examples per task type to prompt the language models. On average, each ReAct example ranges from 352 words to 591 words (590 tokens to 935 tokens). For our study, we reuse the observations, thoughts and actions, and annotate these examples further with the goal and state, which results in a range from 484 to 911 words (807 tokens to 1458 tokens) per example. During our annotation, we discovered minor errors in the ReAct prompts and fixed them as well. In comparison, AdaPlanner Sun et al. (2023), uses a different code-based approach and the prompt has 1104 words (2015 tokens) on average. We use the two-shot examples from ReAct for Alfworld, the one-shot example from ReAct for Webshop and the few-shot prompts of ReAct from ADaPT for Textcraft in all our experiments.